OpenStudy (anonymous):

Find the sum of sin2x and sinx

OpenStudy (anonymous):

I don't understand the question. Don't you just add them together since that's what "sum" means?

OpenStudy (anonymous):

its trig identities

OpenStudy (anonymous):

Do you want it only in terms of sin(x)? Here's a list of identities. http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (across):

The best you can do is simplify the expression:\[ \sin2x+\sin x\\ 2\sin x\cos x+\sin x\\ \sin x(2\cos x+1) \]

OpenStudy (anonymous):

yea in terms of sinx